OpenStudy (anonymous):

find the value of x so that the slope of the line passing through the points (5,3) and (x,6) will be 1 Please help

OpenStudy (tkhunny):

Please calculate the slope between the given points. It will have 'x' in it.

OpenStudy (republican):

Slope is 3/x-5 Since slope is 1 then 1= 3/x-5 x-5 = 3 x = 8 to give a slope of 1

OpenStudy (tkhunny):

3/x-5 is not correct. One should recall one's Order of Operations. (6-3)/(x-5) leads to 3/(x-5) - absolutely NOT the same thing.

OpenStudy (republican):

Disagree ... second pt has coordinates x,6 and first pt has coordinates 3,5. Solving the equation as I showed gives x equal to 8 which means the distance on the x-axis between pt 2 and pt 1 is 3. Therefore slope is 1/1 which is 1.

OpenStudy (anonymous):

no worries i found the answer

OpenStudy (anonymous):

it was 8

OpenStudy (republican):

That's what I showed with the first equation